Regular Meeting 1016
Kokeyama, Sato, Nakano, TTanaka, Yamakoh, Ge
- JPS meeting submission started
- Wednesday is now the DGS maintenance day
- Summary Pages has started
Schedule for Phase-2
Report from the Last 2 weeks
- IMC - ASC: one of the DC PZT was efficiency was too low and swapped. DC loop 40Hz UGF. DAC board was changed to have 2 times gain.
- PSL table
- all lenses are now high power compatible.
- heat sink beam dump will be installed from today.
- PMC temperature loop seems to be working now.
- power cable (banana) to the AOM driver was broken.
- power meters (PDA100) at various positions are installed. One is at the output of the fiber laser output (pickoff). One is the calibrated PMC trans. The last one is the calibrated IMC input.
- IMC MM was improved (TRANS/REFL ratio is 95%).
- Loss may be 15-20% from the fiber laser out to MC TRANS
- Tried 10W for 10 mins. 10W operation has 40Hz harmonics. 20W operation may be advantageous.
- MZM is completely automated.
- Tried to measure the mod index to measure at the PM mon
- 80MHz signal is appearing around through the table and grounds??
- Maybe a rack in the front room of the PSL?
- Beam dumps are being designed.
- OMC installation has started.
- OMMTs installations have started
- Ge is preparing the visa for the next visit, starting in December
- Koji Arai is here!
- Nakano-kun: Nov 1-2
- Kokeyama: Nov 5-
- Yamakoh plan: consult with Miyakawa-san
Nov 12, 10am